OpenStudy (anonymous):

find the parametric representation of the following curve:straight line y=2x+2,z=7x

OpenStudy (loser66):

i think x =t y = 2 + 2t z= 7t

OpenStudy (loser66):

@amistre64 give out the explanation, please

OpenStudy (amistre64):

lol, you answered it, you explain it :)

OpenStudy (anonymous):

how i can find the value of t den?

OpenStudy (loser66):

@amistre64 , please,

OpenStudy (amistre64):

t is the parameter that the others are valued by, hence, parameterixed

OpenStudy (anonymous):

but the answer is like....r(t)=[t,3+2t,7t] how?

OpenStudy (amistre64):

we can write this as: x = 0 + 1x y = 2 +2x z = 0 + 7x let x=y on the right side x = 0 + 1t y = 2 +2t z = 0 + 7t

OpenStudy (amistre64):

r(t) = <x(t),y(t).z(t)> r(t) = <x,y,z> as defined

OpenStudy (amistre64):

 r is a vector from the origin to the point (x,y,z)

OpenStudy (amistre64):

the point (x,y,z) changes as "t" changes x = t y = 2+2t z = 7t r = (t,2+2t,7t)

OpenStudy (anonymous):

how i can write x=0+1x?

OpenStudy (amistre64):

is 0+1x = x a true statement?

OpenStudy (anonymous):

its true but why did u use 0 here?

OpenStudy (anonymous):

wat abot the following qst: ellipse x^2+y^2=1 , z=y pls help me out

OpenStudy (amistre64):

because its a way to express the x component as a sum of the starting point (or vector) and the direction of the x component is heading (the line vector)